# On nodal deformations of singular surfaces in $\mathbb{P}^{3}$

### v1

@@ 1. Introduction
For technical reasons (that we are going to explain in a moment), we have to limit our considerations to regular components of the Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ whose closure may contain the point [S] corresponding to $S$.
@@
Case (a) is considered in Section 3. In Corollary 3.4 we give a general answer to Problem 1.1 for $3\leqslant m\leqslant 5$. This result is proved to be strong for $m=3$ and we may ask if the same is true for $m=4,5$. In Corollary 3.4 we also provide a partial answer for any $m$. In particular we prove that if $S\subset\mathbb{P}^{3}$ is surface of degree $n$ with a unique general ordinary singularity of multiplicity $m$, then the point corresponding to $S$ is in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{{m-1}\choose 2}$. An ordinary multiple point of multiplicity $m$ is said to be general if its tangent cone is the cone over a general plane curve $C$ of degree $m$. In any case, we show that the answer to Problem 1.1 in case (a) definitely depends on the fact that the singularity is ordinary or quasi–ordinary.
@@
The authors also thank Barbara Fantechi and Rita Pardini for useful discussions on related problems to the subject of this paper.
@@ 2. Nodal surfaces with general plane curve section
We will say that $V$ is very regular if for the general $S]\in V$ one has $h^{1}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n-1)\otimes\mathcal{I}_{N})=0$, with $N$ as above the set of nodes of $S$. Of course, if $V$ is very regular, it is also regular. The following proposition is due to Severi in [10, §1]:
@@
(3) $\delta\leqslant{{n+2}\choose 3}-4.$
@@
The regularity hypothesis in Proposition 2.4 is verified for every irreducible component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ if either $n\leqslant 7$ or $n\geqslant 8$ and $\delta\leqslant 4n-5$ (see [8, Thm. 1.1]).
@@
(b) The maximum number of nodes of a sextic surface in $\mathbb{P}^{3}$ is 65 (see, e.g., [9]). A component of the Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(6)|}_{65}$ is regular by Remark 2.5, but is not very regular, because for it the bound in (3) fails. It is a problem to see if there is a very regular, hence good, component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(6)|}_{\delta}$ for $\delta\leqslant 52$. More generally it is not known which is the maximal value of $\delta$ such that there exists a very regular, hence good, component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, with $n\geqslant 6$.
@@ 3. Deformations of surfaces with an ordinary singularity
$t_{n,m}={{n+3}\choose 3}-{{m+2}\choose 3}+2$
and the general point of $\mathcal{T}_{n,m}$ corresponds to a surface with a general ordinary singularity of multiplicity $m$.
@@ 3.3.1. The case of general ordinary multiple point
In order to prove a corollary of this theorem, we need an preliminary lemma.
@@
By applying Proposition 2.6, Remark 2.7, (c), and Lemma 3.3, we have:
@@
    $\mathcal{T}_{n,3}$ sits in the closure of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{4}$ but not in the closure of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, with $\delta>4$;
@@
    $\mathcal{T}_{n,4}$ sits in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{16}$;
@@
    $\mathcal{T}_{n,5}$ sits in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{31}$;
@@
Note that Lemma 3.3 is used only for the second assertion in (i).
@@
Concerning point (i) of Corollary 3.4, notice that $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{4}$ is clearly irreducible and regular. By [8, Thm 1.1], any irreducible component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{16}$ [resp., of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{31}$] is regular as soon as $n\geqslant 6$ [resp., as soon as $n\geqslant 9$].
@@ 3.3.2. The case of a (not necessarily general) ordinary singularity
The invariant $\delta(C)$ can vary with $C$, being upper semi–continuous. This suggests that for a surface $S$ with an ordinary singularity $p$ of multiplicity $m$, the answer to Problem 1.1 could depend on the base curve of the tangent cone to $S$ at $p$. So it is possible that the versal deformation space of an ordinary singularity of multiplicity $m$ is not a topological invariant.
@@ 5.4. Nodal deformations of surfaces with a double line
The proof of Theorem 5.7 shows that $S$, as in the statement of the theorem, can be deformed to a $\delta$–nodal surface (with $\delta=3n-4-\epsilon=2\ell$, where $0\leqslant\epsilon\leqslant 1$) in such a way that $\ell$ (arbitrary chosen) pinch points of $S$ deform each to a pair of nodes.

### v2

@@ 1. Introduction
Our methods are based on deformation theory of regular components of the Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ whose closure may contain the point [S] corresponding to $S$.
@@
Case (a) is considered in Section 3. In Corollary 3.5 we give an answer to Problem 1.1 for $3\leqslant m\leqslant 5$. This is the best possible answer for $m=3$. For $4\leqslant m\leqslant 5$ this is the best possible answer provided one looks only at deformations in regular components of the Severi variety. In Corollary 3.5 we also provide a partial answer for any $m$. In particular we prove that if $S\subset\mathbb{P}^{3}$ is surface of degree $n$ with a unique general ordinary singularity of multiplicity $m$, then the point corresponding to $S$ is in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{{m-1}\choose 2}$. An ordinary singularity of multiplicity $m$ is said to be general if its tangent cone is the cone over a general plane curve $C$ of degree $m$. In any case, we show that the answer to Problem 1.1 in case (a) definitely depends on the fact that the singularity is ordinary or quasi–ordinary.
@@
The authors also thank Fabrizio Catanese, Barbara Fantechi and Rita Pardini for useful discussions on related problems to the subject of this paper.
@@ 2. Nodal surfaces with general plane curve section
We will say that $V$ is very regular if for the general $S]\in V$ one has $h^{1}(\mathbb{P}^{3},\mathcal{O}_{\mathbb{P}^{3}}(n-1)\otimes\mathcal{I}_{N})=0$, with $N$ as above the set of nodes of $S$. Of course, if $V$ is very regular, it is also regular. The following proposition is due to Severi in [11, §1]:
@@
(3) $\delta\leqslant{{n+2}\choose 3}-4=:\delta_{0}(n).$
@@
The regularity hypothesis in Proposition 2.4 is verified for every irreducible component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ if either $n\leqslant 7$ or $n\geqslant 8$ and $\delta\leqslant 4n-5$ (see [9, Thm. 1.1]).
@@
(b) The maximum number of nodes of a sextic surface in $\mathbb{P}^{3}$ is 65 (see, e.g., [10]). A component of the Severi variety $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(6)|}_{65}$ is regular by Remark 2.5, but is not very regular, because for it the bound in (3) fails. It is a problem to see if there is a very regular, hence good, component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(6)|}_{\delta}$ for $\delta\leqslant\delta_{0}(6)=52$. More generally it is not known which is the maximal value of $\delta$ such that there exists a very regular, hence good, component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$, with $n\geqslant 6$. In particular, do there exist, for $n\geqslant 4$, very regular components of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta_{0}(n)}$, where $\delta_{0}(n)$ is defined in (3)?
@@ 3. Deformations of surfaces with an ordinary singularity
$t_{n,m}={{n+3}\choose 3}-{{m+2}\choose 3}+2=\dim(|\mathcal{O}_{\mathbb{P}^{3}}(n)|)-\delta_{0}(m)-1$
(where $\delta_{0}(m)$ is defined in (3)) and the general point of $\mathcal{T}_{n,m}$ corresponds to a surface with a general ordinary singularity of multiplicity $m$.
@@ 3.3.1. The case of general ordinary multiple point
In order to prove a corollary of this theorem, we need some preliminary results.
Proposition 3.3.
Let $(S,p)$ be a germ of surface singularity in $(\mathbb{C}^{3},\bf 0)$, with $p$ an isolated point of multiplicity 3 with tangent cone $\Gamma$ with vertex $p$ over a smooth curve $C$ of degree 3 in $\mathbb{P}^{2}$. Then $(S,p)$ is analitically equivalent to $(\Gamma,p)$.
Proof.
Suppose that $\Gamma$ has equation $f(x,y,z)=0$, where $f(x,y,z)=0$ is a homogeneous polynomial of degree 3, that defines $C$ in $\mathbb{P}^{2}$. By the finite determinacy theorem [7, Thm. 2.23], $(S,p)$ is analitically isomorphic to $(\Gamma,p)$ if
(4) $\mathfrak{m}^{4}\subseteq\mathfrak{m}^{2}\cdot\Big(\frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z}\Big)$
where $\mathfrak{m}$ is the maximal ideal (x,y,z) in $\mathbb{C}[x,y,z]$. Let us prove that (4) holds with in fact the equality. Indeed, consider the multiplication map
$\mu:\mathbb{C}[x,y,z]_{2}\otimes\Big(\frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z}\Big)\longrightarrow\mathbb{C}[x,y,z]_{4},$
where $\mathbb{C}[x,y,z]_{d}$ is the vector space of homogeneous polynomials of degree $d$. Since $C$ is smooth, $\Big(\frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z}\Big)$ is a regular sequence, hence the kernel of $\mu$ has exactly dimension 3, being generated by the Koszul syzygies of $\frac{\partial f}{\partial x},\frac{\partial f}{\partial y},\frac{\partial f}{\partial z}$. Thus
$\dim({\rm Im}(\mu))=3\dim(\mathbb{C}[x,y,z]_{2})-3=15=\dim(\mathbb{C}[x,y,z]_{4})$
that implies the assertion.∎
@@
By applying Proposition 2.6, Remark 2.7, (c), Proposition 3.3 and Lemma 3.4, we have:
@@
    $\mathcal{T}_{n,3}$ sits in the closure of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{4}$ as a subvariety of codimension 3, but not in the closure of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta}$ with $\delta>4$. More precisley, an isolated ordinary triple point of a surface in $\mathbb{P}^{3}$ is never a limit of more than four nodes;
@@
    $\mathcal{T}_{n,4}$ sits in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{16}$, as a codimension $1$ subvariety;
@@
    $\mathcal{T}_{n,5}$ sits in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{31}$, as a codimension $1$ subvariety;
@@
Note that Proposition 3.3 and Lemma 3.4 are used only for (i).
@@
Concerning point (i) of Corollary 3.5, notice that $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{4}$ is clearly irreducible and regular. By [9, Thm 1.1], any irreducible component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{16}$ [resp., of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{31}$] is regular as soon as $n\geqslant 6$ [resp., as soon as $n\geqslant 9$].
@@ 3.3.2. The case of a (not necessarily general) ordinary singularity
If $C$ is general, one has $\delta(C)\leqslant\delta_{0}(m)$. In fact, by Proposition 2.4, a regular and good component of the Severi variety is also very regular, and therefore the assertion follows by Proposition 2.3. As in Remark 2.7, (b), one can ask whether $\delta(C)=\delta_{0}(m)$ if $C$ is general. Suppose that this is the case. Then, by the proof of Theorem 3.7, it would follow that $\mathcal{T}_{n,m}$ sits in the closure of a regular component of $V^{\mathbb{P}^{3},|\mathcal{O}_{\mathbb{P}^{3}}(n)|}_{\delta_{0}(m)}$ as a divisor.
Can the invariant $\delta(C)$ vary with $C$?
@@ 5.4. Nodal deformations of surfaces with a double line
The proof of Theorem 5.7 shows that $S$, as in the statement of the theorem, can be deformed to a $\delta$–nodal surface (with $\delta=3n-4-\epsilon=2\ell$, where $0\leqslant\epsilon\leqslant 1$) in such a way that $\ell$ (arbitrary chosen) pinch points of $S$ deform each to a pair of nodes. This follows by the particular configurations of the nodes of the surface $S_{\Theta}$ (cf. Proposition 4.5 and Lemma 5.1).